OpenStudy (anonymous):

3/x+6/x=1 slove the equation

hero (hero):

Is it: \(\large\frac{3}{x} + \frac{6}{x} = 1\)

OpenStudy (shane_b):

Just multiply everything by x to get rid of the fractions...and the answer will be completely obvious :)

hero (hero):

Why not just add fractions on the right side?

OpenStudy (anonymous):

yes hero

hero (hero):

If you add fractions on the right side, you get: \(\large\frac{9}{x} = 1\) Then you can multiply both sides by x afterwards.

OpenStudy (shane_b):

@Hero: Either way works...I just thought of that way first :)

OpenStudy (anonymous):

1st one to answer wins a cookie :)

hero (hero):

@freysinger, all you need to do is multiply both sides by x. If you can't do that, you need a tutor.

OpenStudy (anonymous):

x=9

OpenStudy (shane_b):

I guess you get a cookie !

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x=9

OpenStudy (shane_b):

@Snapbacklive: You're too late...no cookie for you.

OpenStudy (anonymous):

uhmm man :)